# VQ-CTAP: Cross-Modal Fine-Grained Sequence Representation Learning for Speech Processing

### v1

@@ IV-C ASR Pipeline
VQ-CTAP can be directly applied to the ASR task without fine-tuning, as illustrated in Figure 3(c). In this approach, the weights of the speech encoder and phoneme decoder are frozen. The pre-trained speech encoder is utilized to extract speech embedding $S$. $S$ is quantized using a codebook to obtain the quantized speech embedding $\hat{S}$. Finally, $\hat{S}$ is input into the pre-trained phoneme decoder, which predicts the corresponding phoneme sequence.
@@ V-A Model Details
In the experiment, we use a speech encoder with a structure similar to the Whisper encoder [52], but reduce the dimension to $d$ through a linear layer. The speech encoder consists of 2 convolutional layers, a GELU activation function, 6 transformer layers, and a linear layer. The 2 convolutional layers compress the length to a quarter of the original shape. The phoneme encoder is composed of a convolutional layer, a RELU activation function, 4 transformer layers, and a linear layer. The outputs of the speech encoder and phoneme encoder are layer-normalized separately. The prompt encoder is a VAE-based model, consisting of 6 convolutional layers and an SE-ResNet[53] block. The codebook has 8192 embedding entries, with each embedding having a dimension of 256. The speech decoder comprises 6 transformer layers, 5 convolutional layers, 2 transposed convolutional layers (which increase the length by a factor of 4 to restore the original shape), 6 Tanh activation functions, and a linear layer. The phoneme decoder consists of 6 transformer layers, 2 transposed convolutional layers (which increase the length by a factor of 4 to restore the original length), 2 Tanh activation functions, and a linear layer. Both encoders and decoders have a hidden dimension of 256. The model is trained using 8 NVIDIA TESLA A800 80GB GPUs, with a batch size of 64 per GPU. Adam[54] is used as the optimizer, with an initial learning rate of 2e-4.
@@ V-B Datasets
Regarding the unlabeled speech-only data, we crawl 20,000 hours of speech from the internet, primarily sourced from audiobooks, podcasts, and online videos. However, this portion of the data lacks speaker information statistics. The collected audio underwent various preprocessing steps, including Voice Activity Detection (VAD) segmentation, signal-to-noise ratio (SNR) detection, and human voice detection.
@@ V-C Compared Method and Tasks
To demonstrate the model’s capability for downstream frame-level tasks, we evaluated its performance on TTS, VC, and ASR tasks. We compare our proposed model with five other models, including self-supervised learning methods: Codec[57], Hubert[12], Wav2Vec2.0[9], and supervised learning methods: the encoder of the ASR model Whisper[52], and the speech representation model CTAP [17] based on Contrastive Phoneme-Speech Pretraining. In this paper, the representations extracted from these models are collectively referred to as semantic coding, as shown in Figure1. In all models, the embedding values from the codebook are used to replace discrete codes, resulting in dense representation of the inputs and outputs. To ensure fairness, when constructing the TTS, VC, and ASR systems, only the intermediate semantic coding (representations extracted by these models) is different, the remaining structures are completely identical consistent with Figure3 and trained using the same dataset for fairness. Our model frameworks are aligned with the state-of-the-art methods: SoVITS [^1] https://github.com/svc-develop-team/so-vits-svc for VC and SPEAR-TTS[14] for minimally-supervised TTS. All comparison models were trained from scratch using the same dataset. The memory sizes of the speech encoders for VQ-CTAP, Wav2Vec 2.0, Encodec, HuBERT, and Whisper are 105MB, 310MB, 57MB, 308MB, and 244MB, respectively. While VQ-CTAP has a denser structure compared to other models, it has fewer parameters than HuBERT and Wav2Vec 2.0. Moreover, VQ-CTAP uniquely supports plug-and-play downstream tasks, a feature not offered by the other models.
@@
Figure 5: t-SNE plot of phoneme/speech embedding for 20 speakers
@@ V-D Evaluation Metrics
For the TTS and VC tasks, we conduct all subjective tests using 25 paid human evaluators to assess speech quality, prosody, and timbre similarity.
@@
TABLE II: Results of Minimally-Supervised TTS, VC, and ASR
@@
Figure 7: The Spectrograms, F0, and Energy of synthesized speech (the same semantic coding combining different prompts). The bottom row is the ground-truth.
VI Results
VI-A Compression and Decoupling Ability
Table I compares the compression ratios of VQ-CTAP and mainstream speech representation models. It shows that the 960x compression ratio of VQ-CTAP is 3 times that of Encodec, SoundStream, and Hubert, while the experiments prove that VQ-CTAP maintains high performance(Table II) even at high compression rates.
To intuitively demonstrate that VQ-CTAP’s semantic coding is paralinguistic-independent, we use prompts from 20 speakers to synthesize 4 text, with each synthesis having the same duration to ensure that the text corresponding to each frame of audio is the same across different speakers. We then use VQ-CTAP’s speech encoder to extract the speech representation $S$ and the phoneme encoder to extract the text’s phoneme representation $P$. We expect these two representations to be aligned at the frame level, and $S$ should have removed the speaker information. Using t-SNE, we compress each frame’s $P$ and $S$ to 2-dimensional vectors and show their distribution in Figure 5. The $\bigstar$ represents $P$, and the other shapes represent $S$ for different speakers, with different colors indicating the positions of corresponding $P$ and $S$. As shown in Figures 5(a) and (c), both the Euclidean distance among the $S$ of different speakers and the corresponding $P$ at each frame are very close, proving that VQ-CTAP’s speech encoder extracts the speech representation (semantic coding) while removing paralinguistic information and aligning it with the text modality. Notably, for the red and orange phonemes ”k” in Figure 5(c), although the phonemes are the same and the positions are different, the corresponding $P$ and $S$ are not entangled. This indicates that VQ-CTAP’s representation is highly context-dependent, which is in line with our motivation for using the token-acoustic contrastive loss. The clustering effect in the short sentences (b) and (d) of Figure 5 is not as good as that in the long sentences (a) and (c), also confirming VQ-CTAP’s characteristic of context modeling.
VI-B Evaluation of TTS Task
Table II presents the results of the TTS task. VQ-CTAP achieves the best results in 4 terms of MSEP, WER, P-MOS, and S-MOS. Incorporating the decoder structure and reconstruction loss during training helps extract intermediate representations that are more suitable for frame-level generation tasks. Multiple systems achieve similar results in terms of prosody similarity P-MOS since they employ the same diffusion-based duration model. However, VQ-CTAP achieves the best prosody similarity P-MOS due to its enhanced semantic-paralinguistic consistency, enabled by modules such as the Semantic-Transfer-wise Paralinguistic Consistency Loss. CTAP and VQ-CTAP address the issue of speaker information redundancy in traditional intermediate representations through frame-level contrastive learning. VQ-CTAP exhibits less voice leakage, leading to the best results in terms of voice similarity S-MOS. Furthermore, the pre-trained phoneme encoder alleviates the difficulty of predicting intermediate representations (semantic embeddings) from text in two-stage TTS models and can be directly used as an effective predictor. In terms of speech quality Q-MOS, Codec, Hubert, Wav2Vec2.0, and VQ-CTAP suffer from information loss due to their discrete processing, while CTAP achieves the best Q-MOS results with its continuous representations that are not subject to discretization.
VI-C Evaluation of VC Task
The results of the voice conversion (VC) task are also presented in Table II. Similar to the TTS task, VQ-CTAP achieves the best results in terms of WER, P-MOS, and S-MOS. VQ-CTAP extracts semantic coding from the source speaker’s mel-spectrogram with better decoupling ability for paralinguistic information, and its ability to model the paralinguistic embedding from the target speaker’s mel-spectrogram is also stronger. The experiments reveal that Hubert has a higher information redundancy issue compared to other models, leading to the retention of more paralinguistic information in the semantic encoding, which helps it achieve the best results in terms of Q-MOS. To point out in the ablation study (Table III), CTAP w/o Unsupervised (without using unsupervised low-quality data) achieve similar performance to Hubert in terms of Q-MOS.
VI-D Evaluation of ASR Task
Table II shows that Hubert achieves the best results in ASR, but is not suitable for TTS and VC tasks, because of higher information redundancy. It’s worth noting that in the ablation study (Table III), CTAP w/o Compression (without length compression) achieves results close to Hubert in terms of ACC.
TABLE III: Results of Ablation Studies
@@ VI-D Evaluation of ASR Task
TABLE IV: Frozen pre-training modules comparing independently trained models on the TTS task.
@@
VI-E Ablation Studies
As shown in Table III, the ablation study aims to analyze the contribution of the different components in VQ-CTAP. The largest negative impact of the VQ-CTAP w/o PhonemeDecoder model is on the ACC of the ASR task. The phoneme classification loss of the PhonemeDecoder structure is naturally well-suited for the ASR task. Since contrastive learning has already introduced text supervision, its role is similar to that of the phoneme decoder, so removing the phoneme decoder has little impact on the other results. The VQ-CTAP w/o Compression excels in WER and P-MOS for TTS task and in ACC for ASR task. However, these improvements are relatively minor compared to VQ-CTAP, which maintains excellent reconstruction ability even at high compression rates. Higher compression rates facilitate the separation of paralinguistic information. For autoregressive architectures, such as language models, higher compression rates facilitate easier convergence and enhance stability. We have experimented with compressing HuBERT and wav2vec 2.0 by 2x or 4x, which resulted in significant loss of content information and performance degradation. The VQ-CTAP w/o Contrastive model has the most significant negative impact on all metrics, as the token-acoustic contrastive loss is the source of cross-modal modeling capability. VQ-CTAP w/o Contrastive has degraded to an approximate self-supervised representation learning model, lacking the ability to decouple semantics and paralinguistic information. The VQ-CTAP w/o Consistency model has a decrease in the P-MOS and S-MOS for the TTS and VC tasks. The semantic-transfer-wise paralinguistic consistency loss allows the model to better generalize to unseen data and capture the nuances of paralinguistic information. The VQ-CTAP w/o Unsupervised model has a significant decline in the MSEP, P-MOS, and S-MOS, as the unlabeled data can enhance the representation capability and improve the semantic-paralinguistic decoupling of the representations. Notably, there is an improvement in the Q-MOS metric, due to the removal of the low-quality unlabeled data. The VQ-CTAP w/o Stepping model has a very large negative impact on all metrics. Without the stepping optimization strategy during training, the model’s KL loss and consistency loss are difficult to converge, and the classify loss and MSE loss values increase, leading to a significant decline in the model’s capability. The experiments show that VQ-CTAP has advantages in downstream speech processing tasks, maintaining P-MOS and S-MOS above 3.99 for the TTS and VC tasks.
Table IV verifies the effectiveness of the sequence-aware semantic connector, demonstrating that the method of directly using pre-trained modules for downstream tasks is effective. We separately trained a two-stage TTS model VQ-CTAP + Separate, using the semantic coding extracted by VQ-CTAP, with the text2semantic module and semantic2acoustic module trained separately. The results show that the performance of VQ-CTAP and VQ-CTAP + Separate are very close, indicating that this simple and effective method of using frozen pre-trained modules and adding a connector model is effective.

### v2

@@ IV-C ASR Pipeline
VQ-CTAP can be directly applied to the ASR task without fine-tuning, as illustrated in Figure 3(c). In this approach, the weights of the speech encoder and phoneme decoder are frozen. The pre-trained speech encoder is utilized to extract speech embedding $S$. $S$ is quantized using a codebook to obtain the quantized speech embedding $\hat{S}$. Finally, $\hat{S}$ is fed into the pre-trained phoneme decoder, which predicts the corresponding phoneme sequence.
@@ V-A Model Details
In the experiment, we use a speech encoder with a structure similar to the Whisper encoder [56], but reduce the dimension to $d$ through a linear layer. The speech encoder consists of 2 convolutional layers, a GELU activation function, 6 transformer layers, and a linear layer. The 2 convolutional layers compress the length to a quarter of the original shape. The phoneme encoder is composed of a convolutional layer, a ReLU activation function, 4 transformer layers, and a linear layer. The outputs of the speech encoder and phoneme encoder are layer-normalized separately. The prompt encoder is a VAE-based model, consisting of 6 convolutional layers and an SE-ResNet[57] block. The codebook has 8192 embedding entries, with each embedding having a dimension of 256. The speech decoder comprises 6 transformer layers, 5 convolutional layers, 2 transposed convolutional layers (which increase the length by a factor of 4 to restore the original shape), 6 Tanh activation functions, and a linear layer. The phoneme decoder consists of 6 transformer layers, 2 transposed convolutional layers (which increase the length by a factor of 4 to restore the original length), 2 Tanh activation functions, and a linear layer. Both encoders and decoders have a hidden dimension of 256. The sequence-aware semantic connector consists of 30 residual layers, 64 residual channels, a kernel size of 3, and a dilation cycle of $[1,2,\cdots,512]$. The diffusion step is set to 50, with a linearly spaced schedule $\beta_{t}\in[1\times 10^{-4},0.05]$ ($T=200$). For the stepping optimization strategy, the parameters are as follows: $kl\_start=1e4$, $kl\_end=2e4$, $consistency\_start=2e4$, $consistency\_end=3e4$, $kl\_upper=1e-5$, and $consistency\_upper=1e-5$. The model is trained using 8 NVIDIA TESLA A800 80GB GPUs, with a batch size of 64 per GPU. Adam[58] is used as the optimizer, with an initial learning rate of 2e-4. The number of sample pairs per step ranges from 4,000 to 32,000 during the training process.
@@ V-B Datasets
Regarding the unlabeled speech-only data, we trawl 20,000 hours of speech from the internet, primarily sourced from audiobooks, podcasts, and online videos. However, this portion of the data lacks speaker information statistics. The collected audio underwent various preprocessing steps, including Voice Activity Detection (VAD) segmentation, signal-to-noise ratio (SNR) detection, and human voice detection.
@@ V-C Compared Method and Tasks
To demonstrate the model’s capability for downstream frame-level tasks, we evaluated its performance on TTS, VC, and ASR tasks. We compare our proposed model with five other models, including self-supervised learning methods: Codec[61], Hubert[13], Wav2Vec2.0[10], and supervised learning methods: the encoder of the ASR model Whisper[56], and the speech representation model CTAP [17] based on Contrastive Phoneme-Speech Pretraining. In this paper, the representations extracted from these models are collectively referred to as semantic coding, as shown in TabelII. In all models, the embedding values from the codebook are used to replace discrete codes, resulting in dense representation of the inputs and outputs. To ensure fairness, when constructing the TTS, VC, and ASR systems, only the intermediate semantic coding (representations extracted by these models) is different, the remaining structures are completely identical consistent with Tabe3 and trained using the same dataset for fairness. Our model frameworks are aligned with the state-of-the-art methods: SoVITS [^1] https://github.com/svc-develop-team/so-vits-svc for VC and SPEAR-TTS[9] for minimally-supervised TTS. All comparison models were trained from scratch using the same dataset. The memory sizes of the speech encoders for VQ-CTAP, Wav2Vec 2.0, Encodec, HuBERT, and Whisper are 105MB, 310MB, 57MB, 308MB, and 244MB, respectively. While VQ-CTAP has a denser structure compared to other models, it has fewer parameters than HuBERT and Wav2Vec 2.0. Moreover, VQ-CTAP uniquely supports plug-and-play downstream tasks, a feature not offered by the other models. Many disentangled VQ codecs are multi-codebook models[62, 63, 64], while VQ-CTAP is a single-codebook model. In single-codebook scenarios, VQ-CTAP performs significantly better than these models, which is why we did not choose them as comparison methods.
@@
Figure 5: t-SNE plot of phoneme/speech embedding for 20 speakers. The $\bigstar$ represents $P$, and the other shapes represent $S$ for different speakers, with different colors indicating the positions of corresponding $P$ and $S$. For the red and orange phonemes “k”, although the phonemes are the same and the positions are different, the corresponding $P$ and $S$ are not entangled.
@@ V-D Evaluation Metrics
For the TTS and VC tasks, we conduct all subjective tests using 25 paid human evaluators to assess speech quality, prosody, and timbre similarity. The evaluators have been standardized on a unified 1-5 scoring system, with reference examples provided for each score level. Before participating in the official evaluation, they must pass a test using these standard examples to ensure their understanding and consistency.
@@
TABLE II: Description of Tasks
TTS $text\rightarrow semantic\ prediction\ model\rightarrow$
$intermediate\ representation+prompt\ embedding\rightarrow$
$acoustic\ prediction\ model\rightarrow\ mel$
VC $speech\rightarrow\ semantic\ extractor\rightarrow$
$intermediate\ representation+prompt\ embedding\rightarrow$
$acoustic\ prediction\ model\rightarrow\ mel$
ASR $speech\rightarrow\ semantic\ extractor\rightarrow$
$intermediate\ representation\rightarrow$
$phoneme\ prediction\ model\rightarrow\ phoneme$
TABLE III: Results of Minimally-Supervised TTS, VC, and ASR
@@
Figure 7: The Spectrograms, F0, and Energy of synthesized speech (the same semantic coding combining different prompts). The bottom row is the ground-truth. The synthesized speech and the prompt speech exhibit consistency in the numerical range and variation trends of spectrogram, F0, and energy.
TABLE IV: Results of Ablation Studies
@@
VI Results
VI-A Compression and Decoupling Ability
Table I compares the compression ratios of VQ-CTAP and mainstream speech representation models. It shows that the 960x compression ratio of VQ-CTAP is 3 times that of Encodec, SoundStream, and Hubert, while the experiments prove that VQ-CTAP maintains high performance(Table III) even at high compression rates.
To intuitively demonstrate that VQ-CTAP’s semantic coding is paralinguistic-independent, we use prompts from 20 speakers to synthesize 4 text, with each synthesis having the same duration to ensure that the text corresponding to each frame of audio is the same across different speakers. We then use VQ-CTAP’s speech encoder to extract the speech representation $S$ and the phoneme encoder to extract the text’s phoneme representation $P$. We expect these two representations to be aligned at the frame level, and $S$ should have removed the speaker information. Using t-SNE, we compress each frame’s $P$ and $S$ to 2-dimensional vectors and show their distribution in Figure 5. As shown in Figures 5(a) and (c), both the Euclidean distance among the $S$ of different speakers and the corresponding $P$ at each frame are very close, proving that VQ-CTAP’s speech encoder extracts the speech representation (semantic coding) while removing paralinguistic information and aligning it with the text modality. Notably, for the red and orange phonemes “k” in Figure 5(c), although the phonemes are the same and the positions are different, the corresponding $P$ and $S$ are not entangled. This indicates that VQ-CTAP’s representation is highly context-dependent, which is in line with our motivation for using the token-acoustic contrastive loss. The clustering effect in the short sentences (b) and (d) of Figure 5 is not as good as that in the long sentences (a) and (c), also confirming VQ-CTAP’s characteristic of context modeling.
VI-B Evaluation of TTS Task
Table III presents the results of the TTS task. VQ-CTAP achieves the best results in 4 terms of MSEP, WER, P-MOS, and S-MOS. Incorporating the decoder structure and reconstruction loss during training helps extract intermediate representations that are more suitable for frame-level generation tasks. Multiple systems achieve similar results in terms of prosody similarity P-MOS since they employ the same diffusion-based duration model. However, VQ-CTAP achieves the best prosody similarity P-MOS due to its enhanced semantic-paralinguistic consistency, enabled by modules such as the Semantic-Transfer-wise Paralinguistic Consistency Loss. CTAP and VQ-CTAP address the issue of speaker information redundancy in traditional intermediate representations through frame-level contrastive learning. VQ-CTAP exhibits less voice leakage, leading to the best results in terms of voice similarity S-MOS. Furthermore, the pre-trained phoneme encoder alleviates the difficulty of predicting intermediate representations (semantic embeddings) from text in two-stage TTS models and can be directly used as an effective predictor. In terms of speech quality Q-MOS, Codec, Hubert, Wav2Vec2.0, and VQ-CTAP suffer from information loss due to their discrete processing, while CTAP achieves the best Q-MOS results with its continuous representations that are not subject to discretization.
VI-C Evaluation of VC Task
The results of the voice conversion (VC) task are also presented in Table III. Similar to the TTS task, VQ-CTAP achieves the best results in terms of WER, P-MOS, and S-MOS. VQ-CTAP extracts semantic coding from the source speaker’s mel-spectrogram with better decoupling ability for paralinguistic information, and its ability to model the paralinguistic embedding from the target speaker’s mel-spectrogram is also stronger. The experiments reveal that Hubert has a higher information redundancy issue compared to other models, leading to the retention of more paralinguistic information in the semantic encoding, which helps it achieve the best results in terms of Q-MOS. To point out in the ablation study (Table IV), CTAP w/o Unsupervised (without using unsupervised low-quality data) achieve similar performance to Hubert in terms of Q-MOS.
VI-D Evaluation of ASR Task
Table III shows that Hubert achieves the best results in ASR, but is not suitable for TTS and VC tasks, because of higher information redundancy. It’s worth noting that in the ablation study (Table IV), CTAP w/o Compression (without length compression) achieves results close to Hubert in terms of ACC.
VI-E Ablation Studies
As shown in Table IV, the ablation study aims to analyze the contribution of the different components in VQ-CTAP. The largest negative impact of the VQ-CTAP w/o PhonemeDecoder model is on the ACC of the ASR task. The phoneme classification loss of the PhonemeDecoder structure is naturally well-suited for the ASR task. Since contrastive learning has already introduced text supervision, its role is similar to that of the phoneme decoder, so removing the phoneme decoder has little impact on the other results. The VQ-CTAP w/o Compression excels in WER and P-MOS for TTS task and in ACC for ASR task. However, these improvements are relatively minor compared to VQ-CTAP, which maintains excellent reconstruction ability even at high compression rates. Higher compression rates facilitate the separation of paralinguistic information. For autoregressive architectures, such as language models, higher compression rates facilitate easier convergence and enhance stability. We have experimented with compressing HuBERT and wav2vec 2.0 by 2x or 4x, which resulted in significant loss of content information and performance degradation. The VQ-CTAP w/o Contrastive model has the most significant negative impact on all metrics, as the token-acoustic contrastive loss is the source of cross-modal modeling capability. VQ-CTAP w/o Contrastive has degraded to an approximate self-supervised representation learning model, lacking the ability to decouple semantics and paralinguistic information. The VQ-CTAP w/o Consistency model has a decrease in the P-MOS and S-MOS for the TTS and VC tasks. The semantic-transfer-wise paralinguistic consistency loss allows the model to better generalize to unseen data and capture the nuances of paralinguistic information. The VQ-CTAP w/o Unsupervised model has a significant decline in the MSEP, P-MOS, and S-MOS, as the unlabeled data can enhance the representation capability and improve the semantic-paralinguistic decoupling of the representations. Notably, there is an improvement in the Q-MOS metric, due to the removal of the low-quality unlabeled data. The VQ-CTAP w/o Stepping model has a very large negative impact on all metrics. Without the stepping optimization strategy during training, the model’s KL loss and consistency loss are difficult to converge, and the classify loss and MSE loss values increase, leading to a significant decline in the model’s capability. The experiments show that VQ-CTAP has advantages in downstream speech processing tasks, maintaining P-MOS and S-MOS above 3.99 for the TTS and VC tasks.
Table V verifies the effectiveness of the sequence-aware semantic connector, demonstrating that the method of directly using pre-trained modules for downstream tasks is effective. We separately trained a two-stage TTS model VQ-CTAP + Separate, using the semantic coding extracted by VQ-CTAP, with the text2semantic module and semantic2acoustic module trained separately. The results show that the performance of VQ-CTAP and VQ-CTAP + Separate are very close, indicating that this simple and effective method of using frozen pre-trained modules and adding a connector model is effective.
VI-F Limitations
Regarding dataset selection, our method utilizes a small amount of labeled text-speech pair data alongside a large volume of unlabeled speech-only data. Since the model relies on frame-level alignment information, the quality of the labeled data must be as high as possible. Therefore, we selected the widely recognized open-source datasets: AISHELL3[59] for Chinese and LibriTTS[60] for English. In contrast, the unlabeled data does not require frame-level alignment during training, so we used 20,000 hours of data scraped from the Internet.
Concerning potential biases in the evaluation methodology, existing methods validate semantic representation capabilities using ASR tasks. However, in ASR tasks, the decoupling of paralinguistic information is not strongly correlated with ACC. For example, prosody and emotion can affect recognition results. In the future, we aim to explore better evaluation metrics for semantic representation capabilities.
@@ VI-F Limitations
TABLE V: Frozen pre-training modules comparing independently trained models on the TTS task.